OpenStudy (anonymous):

x/20 = 6.5/0.4 round to the nearest hundredth if necessary

OpenStudy (shamil98):

\[\large \frac{ x }{20} = \frac{ 6.5 }{ 0.4 }\] Use cross multiplication. You result with: \[\large 0.4(x) = 6.5(20)\] Can you do it from here?

OpenStudy (anonymous):

yes I can thank you

OpenStudy (shamil98):

You're welcome.

OpenStudy (anonymous):

okay im just making sure I got it right would it be 325

OpenStudy (shamil98):

Yeah, you're correct.

OpenStudy (anonymous):

ok thanks :)